A portfolio manager purchased 10,000 shares in mutal stock funds and mutual bonds funds valued at $146,000 for a retirement fund. At the time of the purchase, the stock funds sold for $11 per share, while the bond funds sold for $17 per share. How many shares of each did the retirement fund purchase?
Answer by scott8148(6628)

let x=stock, so 10000-x=bond

11(x)+17(10000-x)=146000 ___ 11x+170000-17x=146000 ___ -6x=-24000